# Higher equations of motion at level 2 in Liouville CFT

### v1

@@ 3.2. Expression of the singular state
The first step is a probabilistic expression for the singular state.
@@
We proceed as in Proposition 2.2, except that we are now in a real setting. For all $\varepsilon\in\mathbb{R}$, we introduce the martingale
$\mathcal{E}_{\varepsilon}(t):=e^{e^{2t}(\varepsilon\varphi_{2}-\frac{\varepsilon^{2}}{4}\sinh(2t))},$
@@
By Girsanov’s theorem, the effect of reweighting the measure by $e^{-\varepsilon\varphi_{2}}\mathcal{E}_{\varepsilon}(t)$ is to shift the field $X_{\mathbb{D}}$ as follows
@@
$\quad=e^{\frac{1}{2}(\alpha-Q)c}\frac{\partial}{\partial\varepsilon}_{|\varepsilon=0}\mathbb{E}\left[e^{\frac{\alpha}{2}B_{2t}-\frac{\alpha^{2}}{4}t}\exp\left(-\mu e^{\gamma c}\int_{\mathbb{A}_{t}^{+}}e^{\gamma\frac{\varepsilon}{2}\mathrm{Re}(z^{-2})}\mathrm{d}M_{\gamma}(z)-e^{\frac{\gamma}{2}c}\int_{\mathbb{I}_{t}}e^{\frac{\gamma\varepsilon}{4x^{2}}}\mu_{\partial}(x)\mathrm{d}L_{\gamma}(x)\right)F(\varphi+\frac{\varepsilon}{2}\cos(2\theta))\right]$
@@
$\qquad+\frac{1}{4}\mathbb{E}\left[e^{\frac{\alpha}{2}B_{2t}-\frac{\alpha^{2}}{4}t}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t}^{+})-\mu_{\mathrm{L}}e^{\frac{\gamma}{2}c}L_{\gamma}(\mathbb{I}_{-})-\mu_{\mathrm{R}}e^{\frac{\gamma}{2}c}L_{\gamma}(\mathbb{I}_{+})}\partial_{2}F(\varphi)\right].$
Taking the limit as $t\to\infty$ concludes the proof. ∎
@@ 3.3. First pole of $\mathcal{P}_{\alpha}$ and the $(1,2)$-HEM
It is easy to see that the last line of (3.5) converges and is analytic in a neighbourhood of $\alpha_{1,2}$. Hence, we need only treat the boundary integrals.
We use the estimate $\Psi_{\alpha}^{\partial}(x)=|x|^{-\frac{\gamma\alpha}{2}}(\Psi_{\alpha+\gamma}^{\partial}+O(|x|^{\xi}))$ in $\mathcal{C}^{\prime}$, for some $\xi>0$. This estimate is a straightforward adaptation of Proposition 3.6 in the case of a single $\gamma$-insertion. From this, we get for $\alpha<\alpha_{1,2}$
@@
$=-\frac{2}{\gamma(\alpha-\alpha_{1,2})}\Psi_{\alpha_{-1,2}}^{\partial}+O(1).$
as $\alpha\to\alpha_{1,2}$. The second integral is analytic in the neighbourhood of $\alpha_{1,2}$ since $|x|^{-\frac{\gamma\alpha}{2}-2+\xi}$ is uniformly integrable on $(0,1)$ in the neighbourhood of $\alpha_{1,2}$.
We get the same residue on the other side. Thus, we get $\underset{\alpha=\alpha_{1,2}}{\mathrm{Res}}\,\mathcal{P}_{\alpha}(\varphi_{2})=\frac{1}{2}(\mu_{\mathrm{L}}+\mu_{\mathrm{R}})\Psi_{\alpha_{-1,2}}$ by combining with Proposition 3.2. Finally, Lemma 3.1 allows us to conclude that $\Phi_{\alpha}(\mathbf{S}_{2Q-\alpha}^{0}\mathds{1})$ extends to $\alpha_{1,2}$ with the value
[…]
@@ 3.4. Second pole of $\mathcal{P}_{\alpha}$ and the $(2,1)$-HEM
We have in $e^{-\beta c}\mathcal{H}$
@@
where $\mathcal{R}_{\alpha}$ is analytic at $\alpha_{2,1}$.
@@
Using successively Lemma 3.1, (3.6) and (3.7), we find that $\Phi_{\alpha}(\mathbf{S}_{2Q-\alpha}^{0}\mathds{1})$ extends to $\alpha_{2,1}$ with the value
[…]
$=2\alpha_{2,1}\underset{\alpha=\alpha_{2,1}}{\mathrm{Res}}\left(-\mu\frac{\gamma}{2}(\alpha_{2,1}-\alpha_{1,2})\mathcal{I}_{(2)}(\alpha)+\mu_{\mathrm{L}}^{2}\frac{\gamma^{2}}{8}\mathcal{I}_{(1,1),\eset}^{\partial}(\alpha)-\mu_{\mathrm{R}}\mu_{\mathrm{R}}\frac{\gamma^{2}}{4}\mathcal{I}_{(1),(1)}^{\partial}(\alpha)+\mu_{\mathrm{R}}^{2}\frac{\gamma^{2}}{8}\mathcal{I}_{\eset,(1,1)}(\alpha)\right)$
$=\frac{\gamma^{3}}{8}\left(\mu_{\mathrm{L}}^{2}-2\mu_{\mathrm{L}}\mu_{\mathrm{R}}\cos(\pi\frac{\gamma^{2}}{4})+\mu^{2}_{\mathrm{R}}-\mu\sin(\pi\frac{\gamma^{2}}{4})\right)\frac{\Gamma(\frac{\gamma^{2}}{4})\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(1-\frac{\gamma^{2}}{4})}.$
The remainder of this section is devoted to the proof of Proposition 3.4. The strategy is the same as in the bulk case. To prove (3.7), we use fusion estimates (Proposition 3.6) to factorise the residue into the product of primary field and a Selberg integral. The residue of the Selberg integral can be evaluated explicitly (Lemma 3.5). The analytic continuation of the Poisson operator of (3.6) is handled using integration by parts and a derivative formula for the boundary states.
@@ 3.4.1. Residues of Selberg integrals
By (A.1), the value of the Selberg integral $S_{2,2}(1,-\frac{\gamma\alpha}{2},-\frac{\gamma^{2}}{4})$ is
@@
$\int_{\mathbb{D}\cap\mathbb{H}}|w|^{-\gamma\alpha}|w-\bar{w}|^{-\frac{\gamma^{2}}{2}}\mathrm{Re}(w^{-2})|\mathrm{d}w|^{2}$ $\int_{0}^{1}r^{-\gamma\alpha-\frac{\gamma^{2}}{2}-1}\mathrm{d}r\int_{0}^{\pi}(2\sin\theta)^{-\frac{\gamma^{2}}{2}}\cos(2\theta)\mathrm{d}\theta$
@@
We recall the formula $B(a,b)=2\int_{0}^{\frac{\pi}{2}}\sin(\theta)^{2a-1}\cos(\theta)^{2b-1}\mathrm{d}\theta$ for the Beta function. Using the identity $\Gamma(z+1)=z\Gamma(z)$, the duplication formula $\Gamma(z)\Gamma(z+\frac{1}{2})=\sqrt{\pi}2^{1-2z}\Gamma(2z)$, and the value $\Gamma(\frac{1}{2})=\sqrt{\pi}$, the integral in $\theta$ equals
@@ 3.4.2. Fusion estimates
    Suppose $\alpha+2\gamma<Q$. Then,
    $\Psi_{\alpha}^{\partial}(w)=|w|^{-\gamma\alpha}(2\mathrm{Im}(w))^{-\frac{\gamma^{2}}{2}}(\Psi_{\alpha+2\gamma}^{\partial}+O(|w|^{\xi}));$
    $\Psi_{\alpha}^{\partial}(x_{1},x_{2})=|x_{1}|^{-\frac{\gamma\alpha}{2}}|x_{2}|^{-\frac{\gamma\alpha}{2}}|x_{1}-x_{2}|^{-\frac{\gamma^{2}}{2}}\left(\Psi_{\alpha+2\gamma}^{\partial}+O((|x_{1}|\vee|x_{2}|)^{\xi})\right),$
    for some $\xi>0$.
@@
By item 2. of Proposition 3.6, the last line is absolutely convergent and analytic in the neighbourhood of $\alpha_{2,1}$. By Lemma 3.5, we deduce the following limit in $\mathcal{C}^{\prime}$:
@@
As usual, we can deduce that this equality actually holds in $e^{-\beta c}\mathcal{H}$, which give the first line of (3.7). The proof of the two last lines of (3.7) is identical.
@@
In this case, we have $\alpha_{2,1}+2\gamma>Q$. We only show that $\mathcal{I}_{\eset,(1,1)}$ is regular at $\alpha_{2,1}$. We consider the following regions of $(0,1)^{2}$:
@@
We use the notation $\mathbb{I}_{t}:=(e^{-t},1)$. For $r\in(0,1)$, we have by scaling and item 1. Proposition 3.6
$\int_{r\mathbb{I}_{2}^{2}}\Psi_{\alpha}^{\partial}(x_{1},x_{2})\frac{\mathrm{d}x_{1}}{x_{1}}\frac{\mathrm{d}x_{2}}{x_{2}}=O(r^{-\gamma\alpha-\frac{\gamma^{2}}{2}+\frac{1}{2}(Q-\alpha-2\gamma)^{2}})$
in $\mathcal{C}^{\prime}$. For $\alpha=\alpha_{2,1}$, the exponent equals $\frac{1}{2}(\frac{2}{\gamma}-\gamma)^{2}>0$. Hence, $\int_{\mathcal{D}_{0}}\Psi_{\alpha}^{\partial}(x_{1},x_{2})\frac{\mathrm{d}x_{1}}{x_{1}}\frac{\mathrm{d}x_{2}}{x_{2}}$ is bounded by an absolutely convergent series in the neighbourhood of $\alpha_{2,1}$, so the integral converges and is analytic in this region.
@@ 3.4.3. Meromorphic continuation
This formula is only valid provided all the terms involved are absolutely convergent, which is the case for $\alpha$ in a complex neighbourhood of $\infty$- ∞. It is then extended by analytic continuation to the domain of analyticity of the RHS. Now, we express the analytic continuation of $\mathcal{P}_{\alpha}(\varphi_{2})$ up to $\alpha_{2,1}$. As in the bulk case, the proof relies on a combination of integration by parts and the derivative formula. Fortunately, it happens to be much less tedious.
@@
$(\alpha-\alpha_{1,2})\mathcal{I}_{\eset,(2)}^{\partial}(\alpha)=-\frac{\gamma}{2}\mu_{\mathrm{R}}\mathcal{I}^{\partial}_{\eset,(1,1)}(\alpha)+\frac{\gamma}{2}\mu_{\mathrm{L}}\mathcal{I}_{(1),(1)}(\alpha)+\mathcal{R}_{\alpha};$
@@
where $\mathcal{R}_{\alpha}$, $\mathcal{R}^{\prime}_{\alpha}$ are analytic at $\alpha_{2,1}$.
@@
In the last line, we have defined $\mathcal{R}_{\alpha}$ to be the last two lines of the RHS, which are easily seen to converge and be analytic in a neighbourhood of $\alpha_{2,1}$. For the first two lines, we have symmetrised the singularity $\frac{1}{x_{1}}\frac{1}{x_{2}-x_{1}}$ in order to get the expressions $-\mathcal{I}_{(1),(1)}^{\partial}$ and $\mathcal{I}_{\eset,(1,1)}^{\partial}$. The previous equation is valid for $\alpha<\alpha_{1,2}$, but the RHS is analytic up to $\alpha_{2,1}$, so it expresses the meromorphic extension of $\mathcal{I}_{\eset,(2)}^{\partial}$ in this region. The proof is identical for $\mathcal{I}^{\partial}_{(2),\eset}$.∎
Appendix A Selberg & Dotsenko-Fateev integrals

### v2

@@ 3.2. Expression of the singular state
The first step is to get a probabilistic expression for the singular state.
@@
We proceed as in Proposition 2.2, except that the modes of the field $(\varphi_{n}(t))_{n\;\geqslant\;1}$ under the free field dynamics (3.1) are now real-valued processes. Under this dynamic, these modes evolve as independent Ornstein–Uhlenbeck processes [GRW24b, Section 7.2]. Moreover, the zero mode evolves as an independent Brownian motion of speed $2$: $c_{t}=c+B_{2t}$. For all $\varepsilon\in\mathbb{R}$, we introduce the martingale
$\mathcal{E}_{\varepsilon}(t):=e^{e^{2t}(\varepsilon\varphi_{2}(t)-\frac{\varepsilon^{2}}{4}\sinh(2t))},$
@@
By Cameron–Martin’s theorem, the effect of reweighting the measure by $e^{-\varepsilon\varphi_{2}}\mathcal{E}_{\varepsilon}(t)$ is to shift the field $X_{\mathbb{D}}$ as follows
@@
$\quad=e^{\frac{1}{2}(\alpha-Q)c}\frac{\partial}{\partial\varepsilon}_{|\varepsilon=0}\mathbb{E}\left[e^{\frac{\alpha}{2}B_{2t}-\frac{\alpha^{2}}{4}t}\exp\left(-\mu e^{\gamma c}\int_{\mathbb{A}_{t}^{+}}e^{\gamma\frac{\varepsilon}{2}\mathrm{Re}(z^{-2})}\mathrm{d}M_{\gamma}(z)-e^{\frac{\gamma}{2}c}\int_{\mathbb{I}_{t}}e^{\frac{\gamma\varepsilon}{4x^{2}}}\mu_{\partial}(x)\mathrm{d}L_{\gamma}(x)\right)F(c,\varphi+\frac{\varepsilon}{2}\cos(2\theta))\right]$
@@
$\qquad+\frac{1}{4}\mathbb{E}\left[e^{\frac{\alpha}{2}B_{2t}-\frac{\alpha^{2}}{4}t}e^{-\mu e^{\gamma c}M_{\gamma}(\mathbb{A}_{t}^{+})-\mu_{\mathrm{L}}e^{\frac{\gamma}{2}c}L_{\gamma}(\mathbb{I}_{-})-\mu_{\mathrm{R}}e^{\frac{\gamma}{2}c}L_{\gamma}(\mathbb{I}_{+})}\partial_{2}F\right].$
Taking the limit as $t\to\infty$, the first (resp. second) line of the last equality gives the first (resp. second and third) terms of (3.5). The last line of the last equality converges to $\mathbb{E}[\Psi_{\alpha}^{\partial}\partial_{2}F]$, which has an analytic continutation in a neighbourhood of $(-\infty,Q)$. ∎
@@ 3.3. First pole of $\mathcal{P}_{\alpha}$ and the $(1,2)$-HEM
It is easy to see that the last two terms of (3.5) converge and are analytic in a neighbourhood of $\alpha_{1,2}$. Hence, we only need to treat the boundary integrals (first two terms).
We use the estimate $\Psi_{\alpha}^{\partial}(x)=|x|^{-\frac{\gamma\alpha}{2}}(\Psi_{\alpha+\gamma}^{\partial}+O(|x|^{\xi}))$ in $\mathcal{C}^{\prime}$, for some $\xi>0$ independent of $\alpha$. This estimate is a straightforward adaptation of Proposition 3.7 in the case of a single $\gamma$-insertion. From this, we get the following equality in $\mathcal{C}^{\prime}$, for $\alpha<\alpha_{1,2}$
@@
$=-\frac{2}{\gamma(\alpha-\alpha_{1,2})}\Psi_{\alpha_{1,2}}^{\partial}+O_{\alpha\to\alpha_{1,2}}(1).$
We get the same residue for the term involving the integral over the interval (1,0)(-1,0). Thus, we get $\underset{\alpha=\alpha_{1,2}}{\mathrm{Res}}\,\mathcal{P}_{\alpha}(\varphi_{2})=\frac{1}{2}(\mu_{\mathrm{L}}+\mu_{\mathrm{R}})\Psi_{\alpha_{-1,2}}$ by combining with Proposition 3.3. Finally, Lemma 3.1 allows us to conclude that $\Phi_{\alpha}(\mathbf{S}_{2Q-\alpha}^{0}\mathds{1})$ extends to $\alpha_{1,2}$ with the value
$\Phi_{\alpha_{1,2}}(\mathbf{S}_{2Q-\alpha_{1,2}}^{0}\mathds{1})$ $=\underset{\alpha\to\alpha_{1,2}}{\lim}\,2\alpha(\alpha-\alpha_{2,1})(\alpha-\alpha_{1,2})\mathcal{P}_{\alpha}(\varphi_{2})$
@@ 3.4. Second pole of $\mathcal{P}_{\alpha}$ and the $(2,1)$-HEM
We have in $e^{-\beta c}\mathcal{H}$, for $\alpha<\alpha_{2,1}$,
@@
where $\mathcal{R}_{\alpha}$is analytic up to a neighbourhood of $\alpha_{2,1}$.
@@
The case $\gamma>\sqrt{2}$ is clear from Proposition 3.5 since all residues vanish. Hence, we assume $\gamma<\sqrt{2}$. Using successively Lemma 3.1, (3.6) and (3.7), we find that $\Phi_{\alpha}(\mathbf{S}_{2Q-\alpha}^{0}\mathds{1})$ extends to $\alpha_{2,1}$ with the value
$\Phi_{\alpha_{2,1}}(\mathbf{S}_{2Q-\alpha_{2,1}}^{0}\mathds{1})$ $=\underset{\alpha\to\alpha_{2,1}}{\lim}\,2\alpha(\alpha-\alpha_{1,2})(\alpha-\alpha_{2,1})\mathcal{P}_{\alpha}(\varphi_{2})$
$=2\alpha_{2,1}\underset{\alpha=\alpha_{2,1}}{\mathrm{Res}}\left(-\mu\frac{\gamma}{2}(\alpha_{2,1}-\alpha_{1,2})\mathcal{I}_{(2)}(\alpha)+\mu_{\mathrm{L}}^{2}\frac{\gamma^{2}}{8}\mathcal{I}_{(1,1),\eset}^{\partial}(\alpha)-\mu_{\mathrm{R}}\mu_{\mathrm{R}}\frac{\gamma^{2}}{4}\mathcal{I}_{(1),(1)}^{\partial}(\alpha)+\mu_{\mathrm{R}}^{2}\frac{\gamma^{2}}{8}\mathcal{I}_{\eset,(1,1)}^{\partial}(\alpha)\right)$
$=\frac{\gamma^{2}}{4}\left(\mu_{\mathrm{L}}^{2}-2\mu_{\mathrm{L}}\mu_{\mathrm{R}}\cos(\pi\frac{\gamma^{2}}{4})+\mu^{2}_{\mathrm{R}}-\mu\sin(\pi\frac{\gamma^{2}}{4})\right)\frac{\Gamma(\frac{\gamma^{2}}{4})\Gamma(1-\frac{\gamma^{2}}{2})}{\Gamma(1-\frac{\gamma^{2}}{4})}.$
The remainder of this section is devoted to the proof of Proposition 3.5. The strategy is the same as in the bulk case. To prove (3.7), we use fusion estimates (Proposition 3.7) to factorise the residue into the product of a primary field and a Selberg integral. The residue of the Selberg integral can be evaluated explicitly (Lemma 3.6). The analytic continuation of the Poisson operator of (3.6) is handled using integration by parts and a derivative formula for the boundary states.
@@ 3.4.1. Residues of Selberg integrals
We refer to Appendix A for the definition of the Selberg integrals $S_{2,2}$ and $S_{2,1}$. By (A.1), the value of $S_{2,2}(1,-\frac{\gamma\alpha}{2},-\frac{\gamma^{2}}{4})$ is
@@
$\int_{\mathbb{D}\cap\mathbb{H}}|w|^{-\gamma\alpha}|w-\bar{w}|^{-\frac{\gamma^{2}}{2}}\mathrm{Re}(w^{-2})|\mathrm{d}w|^{2}$ $=\int_{0}^{1}r^{-\gamma\alpha-\frac{\gamma^{2}}{2}-1}\mathrm{d}r\int_{0}^{\pi}(2\sin\theta)^{-\frac{\gamma^{2}}{2}}\cos(2\theta)\mathrm{d}\theta$
@@
We recall the formula $B(a,b)=2\int_{0}^{\frac{\pi}{2}}\sin(\theta)^{2a-1}\cos(\theta)^{2b-1}\mathrm{d}\theta$ for the Beta function, and the expression $B(a,b)=\frac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}$. Using the identity $\Gamma(z+1)=z\Gamma(z)$, Euler’s formula $\Gamma(z)\Gamma(1-z)=\frac{\pi}{\sin(\pi z)}$, the duplication formula $\Gamma(z)\Gamma(z+\frac{1}{2})=\sqrt{\pi}2^{1-2z}\Gamma(2z)$, and the value $\Gamma(\frac{1}{2})=\sqrt{\pi}$, the integral in $\theta$ equals
@@ 3.4.2. Fusion estimates
    Suppose $\alpha+2\gamma<Q$ and define $\Delta:=\{(x_{1},x_{2})\in(\mathbb{I}\setminus\{0\})^{2}|\,|\log|x_{2}/x_{1}||\;\geqslant\;1\}$. There exists $\xi>0$ such that:
    $\Psi_{\alpha}^{\partial}(w)=|w|^{-\gamma\alpha}(2\mathrm{Im}(w))^{-\frac{\gamma^{2}}{2}}(\Psi_{\alpha+2\gamma}^{\partial}+O(|w|^{\xi})),$ $\text{uniformly over }\mathbb{D}\cap\mathbb{H};$
    $\Psi_{\alpha}^{\partial}(x_{1},x_{2})=|x_{1}|^{-\frac{\gamma\alpha}{2}}|x_{2}|^{-\frac{\gamma\alpha}{2}}|x_{1}-x_{2}|^{-\frac{\gamma^{2}}{2}}\left(\Psi_{\alpha+2\gamma}^{\partial}+O((|x_{1}|\vee|x_{2}|)^{\xi})\right),$ $\text{uniformly over }\Delta.$
@@
By item 2. of Proposition 3.7, the last line is absolutely convergent and analytic in a neighbourhood of $\alpha_{2,1}$. By Lemma 3.6, we deduce the following limit in $\mathcal{C}^{\prime}$:
@@
Reproducing the argument at the end of the proof of Proposition 2.3, we deduce that this equality actually holds in $e^{-\beta c}\mathcal{H}$, which give the first line of (3.7). The proof of the last two lines of (3.7) is identical, with the only difference that we must first remove the region where $\log|x_{2}/x_{1}|<1$ where the fusion estimate of the second item of Proposition 3.7 does not hold. This strategy has already been carried out in the proof of (2.8), and we omit the details.
@@
In this case, we have $\alpha_{2,1}+2\gamma>Q$. We only show that $\mathcal{I}_{\eset,(1,1)}$ is regular at $\alpha_{2,1}$, the other case being dealt with similarly. We consider the following regions of $(0,1)^{2}$:
@@
We use the notation $\mathbb{I}_{t}:=(e^{-t},1)$. For $r\in(0,1)$, we have by scaling and the first item of Proposition 3.7
$\int_{r\mathbb{I}_{2}^{2}}\Psi_{\alpha}^{\partial}(x_{1},x_{2})\frac{\mathrm{d}x_{1}}{x_{1}}\frac{\mathrm{d}x_{2}}{x_{2}}=O(r^{-\gamma\alpha-\frac{\gamma^{2}}{2}+\frac{1}{2}(Q-\alpha-2\gamma)^{2}})\qquad\text{in }\mathcal{C}^{\prime}$
For $\alpha=\alpha_{2,1}$, the exponent equals $\frac{1}{2}(\frac{2}{\gamma}-\gamma)^{2}>0$. Hence, $\int_{\mathcal{D}_{0}}\Psi_{\alpha}^{\partial}(x_{1},x_{2})\frac{\mathrm{d}x_{1}}{x_{1}}\frac{\mathrm{d}x_{2}}{x_{2}}$ is bounded by an absolutely convergent series in a neighbourhood of $\alpha_{2,1}$, so the integral converges and is analytic in this region.
@@ 3.4.3. Meromorphic continuation
This formula is only valid provided all the terms involved are absolutely convergent, which is the case for $\alpha$ in a complex neighbourhood of -$\infty$. It is then extended by analytic continuation to the domain of analyticity of the RHS.
Now, we express the analytic continuation of $\mathcal{P}_{\alpha}(\varphi_{2})$ up to $\alpha_{2,1}$. As in the bulk case, the proof relies on a combination of integration by parts and the derivative formula. Fortunately, it happens to be much less tedious.
@@
$(\alpha-\alpha_{1,2})\mathcal{I}_{\eset,(2)}^{\partial}(\alpha)=-\frac{\gamma}{2}\mu_{\mathrm{R}}\mathcal{I}^{\partial}_{\eset,(1,1)}(\alpha)+\frac{\gamma}{2}\mu_{\mathrm{L}}\mathcal{I}_{(1),(1)}^{\partial}(\alpha)+\mathcal{R}_{\alpha};$
@@
where $\mathcal{R}_{\alpha}$, $\mathcal{R}^{\prime}_{\alpha}$ are regular at $\alpha_{2,1}$.
@@
In the last line, we have defined $\mathcal{R}_{\alpha}$ to be the last two lines of the RHS, which are easily seen to converge and be analytic in a neighbourhood of $\alpha_{2,1}$. For the first two lines, we have symmetrised the singularity $\frac{1}{x_{1}}\frac{1}{x_{2}-x_{1}}$ in order to get the expressions $\mathcal{I}_{(1),(1)}^{\partial}$ and $\mathcal{I}_{\eset,(1,1)}^{\partial}$. The previous equation is valid for $\alpha<\alpha_{1,2}$, but the RHS is analytic up to $\alpha_{2,1}$, so it expresses the meromorphic extension of $\mathcal{I}_{\eset,(2)}^{\partial}$ in this region. The proof is identical for $\mathcal{I}^{\partial}_{(2),\eset}$.∎
Appendix A Selberg & Dotsenko–Fateev integrals